OpenStudy (anonymous):

what is the graph of the function f(x) = -x^2+x+20/x+4

OpenStudy (ybarrap):

f(x)= (-x^2+x+20)/(x+4) But -x^2+x+20 = (5-x) × (x+4) So, (-x^2+x+20)/(x+4) = -((x+4) (x-5)) / (x + 4)= 5 - x y = -x + 5 Graph intercepts the y-axis at y = 5 and has a slope of -1. Make sense?